### v1

@@ V-B Experimental Setup
V-C The Impact of AoI on Inference Error
Fig. 5 illustrates how the inference error varies with the AoI for two modalities, with each AoI ranging from 1 to 50. The color intensity or surface height represents the expected inference error on the testing dataset. Although the inference error generally increases with the AoI of either modality, the function is not monotonic. Furthermore, the impact of each modality differs significantly: as the AoI of modality 1 (LiDAR) increases, the inference error grows significantly faster than it does for modality 2 (control signal). It indicates that LiDAR data is more strongly correlated with the target signal. Moreover, the figure suggests that the inference error may not be an additive function of the AoI vector; the effect of one modality’s AoI on the inference error depends on the AoI of the other modality. Overall, the empirical results show that the inference error may be a non-monotonic and non-additive function of the AoI vector.
V-D Scheduling Policies Evaluation
@@
    Index-based threshold policy: This is our policy, as described in Theorem 1.
@@
    Uniform random policy: The policy randomly selects one of the two modalities with equal probability, regardless of the current AoI state.
We use the results obtained in Section V-C as the empirical expected inference error function of the AoI vector. We vary the transmission time for each modality, taking values $2,4,6,8,10$, to reflect different feature sizes. Fig. 5 presents the performance of each scheduling policy under varying transmission times. Our proposed index policy consistently achieves the best performance across all cases, reducing the inference error by up to 53% compared to the randomized policy and up to 55% compared to the round-robin policy. Although the round-robin policy is effective for minimizing AoI, it fails to capture the complex relationship between inference error and AoI, leading to suboptimal performance in certain scenarios (e.g., when $T_{1}=2$ and $T_{2}=10$). The uniform random policy exhibits similar limitations, as it ignores both inference error and AoI. These results underscore the importance of jointly considering both modalities and their impact on inference error when designing scheduling policies.
VI Conclusion and future work
We studied the two-modality scheduling problem for remote inference systems, where a receiver-side ML model relies on time-sensitive data from remote sensors. We derived an optimal policy applicable to general AoI penalty functions and heterogeneous transmission times. Experiments on robot state prediction show that our policy reduces inference error by up to 55% compared to baselines.
One future direction is to extend the proposed index-based threshold policy beyond two coupled modalities to general multimodal settings. The extension is non-trivial due to the curse of dimensionality, as the state space grows exponentially with more modalities. Another direction is to relax the current assumptions of reliable communication and deterministic transmission times. Considering unreliable channels and random delays would make the model more practical, but also add significant complexity to policy design. Last but not least, we assume that ML inference error is a time-invariant AoI function. Optimizing a time-dependent AoI function is essential for applications with complex, non-stationary dynamics.

### v2

@@ V-A Experimental Setup
V-B The Impact of AoI on Inference Error
Fig. 5 illustrates how the inference error varies with the AoI for two modalities, with each AoI ranging from 1 to 50. The color intensity or surface height represents the expected inference error on the testing dataset. Although the inference error generally increases with the AoI of either modality, the function is not monotonic. Furthermore, the impact of each modality differs significantly: as the AoI of modality 1 (LiDAR) increases, the inference error grows faster than it does for modality 2 (control signal). It indicates that LiDAR data is more strongly correlated with the target signal. Moreover, the figure suggests that the inference error may not be an additive function of the AoI vector; the effect of one modality’s AoI on the inference error depends on the AoI of the other modality. Overall, the results show that the inference error may be a non-monotonic and non-additive function of the AoI vector.
V-C Scheduling Policies Evaluation
@@
    Index-based threshold policy: This is our policy described in Theorem 1. Note that the index function is only needed to determine the optimal solution $(\tau_{1}^{*},\tau_{2}^{*})$; once determined, the scheduler only uses $(\tau_{1}^{*},\tau_{2}^{*})$ during operation.
@@
    Uniform random policy: The policy randomly selects one of the two modalities with equal probability.
We use the results obtained in Section V-B as the empirical expected inference error function. We vary the transmission time for each modality, taking values $2,4,6,8,10$, to reflect different feature sizes. Fig. 5 presents the performance of each scheduling policy under varying transmission times. Our proposed index policy consistently achieves the best performance across all cases. Particularly, when $T_{1}$ = 2 and $T_{2}$ = 6, our policy reduces the inference error by 55% compared with the round-robin policy and by 53% compared with the uniform random policy. Note that while the round-robin policy is effective for minimizing AoI [15], it fails to capture the complex relationship between inference error and AoI, leading to suboptimal performance in certain scenarios (e.g., when $T_{1}=2$ and $T_{2}=10$). The uniform random policy exhibits similar limitations. These results underscore the importance of jointly considering both modalities and their impact on inference error when designing scheduling policies.
VI Conclusion
We studied the two-modality scheduling problem for remote inference systems, where a receiver-side ML model relies on time-sensitive data from remote sensors. We developed an optimal policy applicable to general AoI penalty functions and heterogeneous transmission times. Experiments on robot state prediction show that our policy reduces inference error by up to 55% compared with baselines. One future extension is to study scheduling over unreliable channels.